the A swimming pool is the shape of a rectangle. the length of the pool is 12 feet more than the width. A concrete deck 3 feet wide is added around the pool, and perimeter of the deck is 148 feet. Find the dimensions of the swimming pool.
Answer by macston(5194)

The walk adds 3 feet to each end of each side , so (2 ends)(4 sides)(3ft)=24 feet The perimeter of the walk is 24 feet more than the perimeter of the pool. The perimeter of the pool is 148 feet-24 feet=124 feet.
P=perimeter; W=width; L=length=W+12ft
P=2(W+L) Substitute for L.
124ft=2(W+(W+12ft)) Divide each side by 2.
62ft=2W+12ft Subtract 12ft from each side.
50ft=2W divide each side by 2.
25ft=W ANSWER 1: The width of the pool is 25 feet.
L=W+12ft=25ft+12ft=37ft ANSWER 2: The length of the pool is 37 feet.